# Lie algebras generated by reflections in types BCD

### v1

@@ 5.1. Statement of the main theorem
$0\textstyle 2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=(2n-8)+\binom{n}{i}\cdot(n-3-i)+2\cdot\binom{n-1}{i-1}>$ 0 .
On the other hand, for $i=n-2$ one has
@@
Then $02\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=\binom{n}{2}-4>$ 0. Thus, if $(\lambda,\mu)=([n-1-i,1],[1^{i}])$ for some $21\leq i\leq n-$ 2, then $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{(\lambda,\mu)})$, as desired.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (5) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The right side of (5.5.1) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by the induction hypothesis that (5.5.1) is a surjection. By induction, one has
@@
$0\textstyle 2\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq(\frac{n}{2}-1)\cdot\binom{n-2}{m-1}+(2n-6)>$ 0 .
Then $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is a simple Lie algebra by [13, Lemme 15]. Since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, the preceding inequalities also imply that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$. Then since $n\geq 6$, it follows from [13, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
The right side of (5.5.2) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by the induction hypothesis that (5.5.2) is a surjection. By induction, one has
@@
Since $\mathfrak{d}_{\left\{\lambda,[1]\right\}}\subseteq\mathfrak{osp}\{\lambda,[1]\}$, then $\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})\leq\frac{1}{2}\cdot\dim(S^{\left\{\lambda,[1]\right\}})=\frac{n}{2}\cdot f^{\lambda}$. Also, since $\lambda=\lambda^{*}$ and $n-1\geq 5$, then $f^{\lambda}\geq(n-1)-1\geq 2$ by [10, Theorem 2.4.10]. Then
@@
so $\mathfrak{d}_{\left\{\lambda,[1]\right\}}$ is simple by [13, Lemme 15]. Further, $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})\leq\dim(S^{\left\{\lambda,[1]\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})$, and then one deduces using [13, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,[1]\right\}}=\mathfrak{osp}\{\lambda,[1]\}$.
@@
where the $\star$ indicates that the term $\mathfrak{r}_{\left\{\mu,-\right\}}$ is omitted if $|\mu|$ is odd. Then the right side of (5.5.3) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by the induction hypothesis that (5.5.3) is a surjection. By induction, one has
@@
where at the second inequality we note that the additional $1$ is subtracted because there may not be any $\tau\prec\mu$ such that $\operatorname{res}(\mu/\tau)=0$ (and hence the $-2$ term on the first line results in subtracting one more than $r_{\lambda}+r_{\mu}$). Since $\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{osp}\{\lambda,\mu\}$, then $\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\frac{1}{2}\cdot\dim(S^{\left\{\lambda,\mu\right\}})$, and hence
@@
Then $02\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq 2n-(n+1)=n-1>$ 0, so $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is simple by [13, Lemme 15]. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, and then deduces using [13, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
The right side of (5.5.4) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by the induction hypothesis that (5.5.4) is a surjection. By induction, one has
@@
Then $02\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq n>$ 0, so $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is simple by [13, Lemme 15]. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, and then deduces using [13, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (6) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The right side of (5.6.1) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by the induction hypothesis that (5.6.1) is a surjection.
@@
This implies by [13, Lemme 15] that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ is a simple Lie algebra. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, and then deduces using [13, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$. ∎
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (7) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
Now if $2\cdot\operatorname{rk}(\mathfrak{g})>\dim(V)$, then $\mathfrak{g}\cong\mathfrak{sl}(V)$ by [13, Lemme 13], as desired. Otherwise one has
@@
One has $\dim(V)\geq 2\cdot\binom{n}{n/2}$ as in (5.3.2), and since $n\geq 6$, this implies that $\dim(V)\geq 40$. Then by [13, Lemma 14], exactly one of the following must be true:
@@
[…]
@@
[…]
@@
Let $R=R^{(\lambda,\lambda)}\in\mathbb{T}(\lambda,\lambda)_{+1}$ be the row major bitableau defined prior to Lemma 2.4.4. Rescaling the form (,)(-,-)( - , - ) if necessary, we may assume that $(u_{R^{*}},u_{R})=1$. Let $T\in\mathbb{T}(\lambda,\lambda)_{+1}$. The proof of [15, Lemma 6.2] shows that $T$ can be obtained from $R$ by applying a sequence of $\ell(T)$ admissible adjacent transpositions, none of which move the integer $1$, so at each intermediate step one still has an element of $\mathbb{T}(\lambda,\lambda)_{+1}$. Then using (5.7.1), one can show for arbitrary $T\in\mathbb{T}(\lambda,\lambda)_{+1}$ that
@@
    for some $21\leq i\leq n-$ 2, and $\left\{\nu,\tau\right\}$ is the unique arm and leg such that $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (8) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
where the $\star$ indicates that the term $\mathfrak{r}_{\left\{\nu,-\right\}}$ is omitted if $\nu=\nu^{*}$ and $|\nu|$ is odd. It follows from Lemma 5.8.1 that the right side of (5.8.1) is a summand of the right side of (4.3.16) for the value $n-1$, except perhaps that a summand of the form $\mathfrak{r}_{\beta_{n-1,1}}$ may be replaced by an isomorphic summand of the form $\mathfrak{r}_{\beta_{n-1,d}}$ for some $21\leq d\leq n-$ 2, and a summand of the form $\mathfrak{s}_{\alpha_{n-1,1}}$ may be replaced by an isomorphic summand $\mathfrak{s}_{\alpha_{n-1,d}}$ for some $31\leq d\leq n-$ 3. This implies by induction that (5.8.1) is a surjection, and implies for each $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$ with $\nu\neq\tau$ (resp. $\nu=\tau$) that $\mathfrak{r}_{\left\{\nu,\tau\right\}}$ is a simple Lie algebra acting irreducibly on $S^{\left\{\nu,\tau\right\}}$ (resp. that $\mathfrak{r}_{\left\{\nu,\pm\right\}}$ is a simple acting irreducibly on $S^{\left\{\nu,\pm\right\}}$).
Let $\mathfrak{h}=\rho_{\left\{\lambda,\mu\right\}}(\mathfrak{r}^{\prime})$. If $\left\{\nu,\nu\right\}\prec\left\{\lambda,\mu\right\}$ with $\nu=\nu^{*}$ and $|\nu|$ odd, then $\mathfrak{r}_{\left\{\nu,\pm\right\}}\cong\mathfrak{sl}\{\nu,\pm\}$ by induction, and it follows that $S^{\left\{\nu,+\right\}}$ and $S^{\left\{\nu,-\right\}}$ are dual but not isomorphic as $\mathfrak{h}$-modules because $\dim(S^{\left\{\nu,\pm\right\}})=\frac{1}{2}\cdot\binom{n-1}{|\nu|}\cdot f^{\nu}\cdot f^{\nu}\geq\frac{1}{2}\cdot(n-1)>2$, and the natural module for $\mathfrak{sl}(k)$ is self-dual only if $k=2$. Now $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ such that the restriction of $S^{\left\{\lambda,\mu\right\}}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. Since $\mathfrak{h}\subseteq\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{sl}\{\lambda,\mu\}$, then $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})$. Our goal now is to show that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{\left\{\lambda,\mu\right\}})$. This will imply first by [13, Lemme 15] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is a simple Lie algebra, and then by [13, Lemme 13] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{sl}\{\lambda,\mu\}$. We break the remainder of the proof into the following cases:
@@
If $\tau\prec\lambda$, then $\dim(S^{\left\{\tau,[1]\right\}})=(n-1)\cdot f^{\tau}\geq 5$. And since $\left\{\lambda,[1]\right\}$ is not an arm and leg, then $\lambda$ is not either $[n-1]$ or $[1^{n-1}]$, and hence $\dim(S^{\left\{\lambda,\emptyset\right\}})=f^{\lambda}\geq(n-1)-1\geq 4$. This implies that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{\left\{\lambda,[1]\right\}})$, as desired.
@@
Suppose there exists an arm and leg $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$, and all $\left\{\eta,\sigma\right\}\prec\left\{\lambda,\mu\right\}$ are proper. Replacing $\left\{\lambda,\mu\right\}$ with $\left\{\lambda^{*},\mu^{*}\right\}$ if necessary, we may assume that $\left\{\lambda,\mu\right\}=\left\{[n-1-i,1],[1^{i}]\right\}$ for some $22\leq i\leq n-$ 2. If $\left\{\nu,\tau\right\}\neq\left\{\nu^{*},\tau^{*}\right\}$ for all $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$, then $02\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\mu\right\}})>$ 0 as in Section 5.4.4, so suppose $\left\{\nu,\tau\right\}=\left\{\nu^{*},\tau^{*}\right\}$ for some $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$. Then $n=2i+1$, $\left\{\lambda,\mu\right\}=\left\{[i,1],[1^{i}]\right\}$, and as $\mathfrak{r}$-modules,
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value $n-1$. Then part (9) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.9.1 that the right side of (5.9.1) is a summand of the right side of (4.3.17) for the value $n-1$. This implies by induction that (5.9.1) is a surjection, and implies for each $\nu\prec\lambda$ that $\mathfrak{r}_{\left\{\nu,\lambda\right\}}\cong\mathfrak{sl}\{\nu,\lambda\}$. Let $\mathfrak{h}=\rho_{\left\{\lambda,\pm\right\}}(\mathfrak{r}^{\prime})$. Then $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ such that the restriction of $S^{\left\{\lambda,\pm\right\}}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. One has

### v2

@@ 5.1. Statement of the main theorem
$2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=(2n-8)+\binom{n}{i}\cdot(n-3-i)+2\cdot\binom{n-1}{i-1}>0.$
On the other hand, for i=n2i=n-2 one has
@@
Then $2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{(\lambda,\mu)})=\binom{n}{2}-4>0$. Thus, if $(\lambda,\mu)=([n-1-i,1],[1^{i}])$ for some $1\leq i\leq n-2$, then $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{(\lambda,\mu)})$, as desired.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (5) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The right side of (5.5.1) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by the induction hypothesis that (5.5.1) is a surjection. By induction, one has
@@
$2\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq(\frac{n}{2}-1)\cdot\binom{n-2}{m-1}+(2n-6)>0.$
Then $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is a simple Lie algebra by [Marin:2007, Lemme 15]. Since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, the preceding inequalities also imply that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$. Then since $n\geq 6$, it follows from [Marin:2007, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
The right side of (5.5.2) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by the induction hypothesis that (5.5.2) is a surjection. By induction, one has
@@
Since $\mathfrak{d}_{\left\{\lambda,[1]\right\}}\subseteq\mathfrak{osp}\{\lambda,[1]\}$, then $\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})\leq\frac{1}{2}\cdot\dim(S^{\left\{\lambda,[1]\right\}})=\frac{n}{2}\cdot f^{\lambda}$. Also, since $\lambda=\lambda^{*}$ and $n-1\geq 5$, then $f^{\lambda}\geq(n-1)-1\geq 2$ by [James:1981, Theorem 2.4.10]. Then
@@
so $\mathfrak{d}_{\left\{\lambda,[1]\right\}}$ is simple by [Marin:2007, Lemme 15]. Further, $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})\leq\dim(S^{\left\{\lambda,[1]\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,[1]\right\}})$, and then one deduces using [Marin:2007, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,[1]\right\}}=\mathfrak{osp}\{\lambda,[1]\}$.
@@
where the $\star$ indicates that the term $\mathfrak{r}_{\left\{\mu,-\right\}}$ is omitted if $|\mu|$ is odd. Then the right side of (5.5.3) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by the induction hypothesis that (5.5.3) is a surjection. By induction, one has
@@
where at the second inequality we note that the additional $1$ is subtracted because there may not be any $\tau\prec\mu$ such that $\operatorname{res}(\mu/\tau)=0$ (and hence the 2-2 term on the first line results in subtracting one more than $r_{\lambda}+r_{\mu}$). Since $\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{osp}\{\lambda,\mu\}$, then $\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\frac{1}{2}\cdot\dim(S^{\left\{\lambda,\mu\right\}})$, and hence
@@
Then $2\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq 2n-(n+1)=n-1>0$, so $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is simple by [Marin:2007, Lemme 15]. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, and then deduces using [Marin:2007, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
The right side of (5.5.4) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by the induction hypothesis that (5.5.4) is a surjection. By induction, one has
@@
Then $2\cdot\operatorname{rk}(\mathfrak{h})-\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\geq n>0$, so $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is simple by [Marin:2007, Lemme 15]. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})$, and then deduces using [Marin:2007, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{osp}\{\lambda,\mu\}$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (6) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
The right side of (5.6.1) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by the induction hypothesis that (5.6.1) is a surjection.
@@
This implies by [Marin:2007, Lemme 15] that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ is a simple Lie algebra. Finally, since $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\pm\right\}})$, one sees that $2\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\pm\right\}})\leq\dim(S^{\left\{\lambda,\pm\right\}})<4\cdot\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\pm\right\}})$, and then deduces using [Marin:2007, Lemme 14] that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}=\mathfrak{osp}\{\lambda,\pm\}$. ∎
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (7) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
Now if $2\cdot\operatorname{rk}(\mathfrak{g})>\dim(V)$, then $\mathfrak{g}\cong\mathfrak{sl}(V)$ by [Marin:2007, Lemme 13], as desired. Otherwise one has
@@
One has $\dim(V)\geq 2\cdot\binom{n}{n/2}$ as in (5.3.2), and since $n\geq 6$, this implies that $\dim(V)\geq 40$. Then by [Marin:2007, Lemma 14], exactly one of the following must be true:
@@
Suppose there exists a nondegenerate $\mathfrak{g}$-invariant symmetric bilinear form (,)(-,-) on $V$. Thus for all $u,v\in V$ and $x\in\mathfrak{g}$, one has $(u,v)=(v,u)$ and (x.u,v)=(u,x.v)(x.u,v)=-(u,x.v). By abuse of notation, we will conflate an element $x\in\mathfrak{d}_{n}^{\prime}$ with its image in $\mathfrak{g}=\mathfrak{d}_{\left\{\lambda,+\right\}}$. Let $p:\mathfrak{d}_{n}\to Z(\mathfrak{d}_{n})$ be the projection map of Lemma 4.1.1. One has $Z(\mathfrak{d}_{n})=\operatorname{span}\{\mathcal{X}_{n}\}$ by Lemma 4.2.3, and since $\lambda=\lambda^{*}$, one sees that $\mathcal{X}_{n}$ acts as zero on $S^{\left\{\lambda,\lambda\right\}}$, and hence as zero on $V$. Then for arbitrary $z\in\mathfrak{d}_{n}$, one has
@@
and hence (,)(-,-) is $\mathfrak{d}_{n}$-invariant. This implies for $s$ in the set $\left\{s_{1},\ldots,s_{n-1},\widetilde{s}_{n}\right\}$ of generators for $\mathfrak{d}_{n}$ that (s.u,s.v)=(u,v)(s.u,s.v)=-(u,v). Since these elements generate $\mathcal{D}_{n}$, it follows for $g\in\mathcal{D}_{n}$ that
@@
Let $R=R^{(\lambda,\lambda)}\in\mathbb{T}(\lambda,\lambda)_{+1}$ be the row major bitableau defined prior to Lemma 2.4.4. Rescaling the form (,)(-,-) if necessary, we may assume that $(u_{R^{*}},u_{R})=1$. Let $T\in\mathbb{T}(\lambda,\lambda)_{+1}$. The proof of [Mishra:2016, Lemma 6.2] shows that $T$ can be obtained from $R$ by applying a sequence of $\ell(T)$ admissible adjacent transpositions, none of which move the integer $1$, so at each intermediate step one still has an element of $\mathbb{T}(\lambda,\lambda)_{+1}$. Then using (5.7.1), one can show for arbitrary $T\in\mathbb{T}(\lambda,\lambda)_{+1}$ that
@@
    for some $1\leq i\leq n-2$, and $\left\{\nu,\tau\right\}$ is the unique arm and leg such that $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$.
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (8) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
where the $\star$ indicates that the term $\mathfrak{r}_{\left\{\nu,-\right\}}$ is omitted if $\nu=\nu^{*}$ and $|\nu|$ is odd. It follows from Lemma 5.8.1 that the right side of (5.8.1) is a summand of the right side of (4.3.16) for the value n1n-1, except perhaps that a summand of the form $\mathfrak{r}_{\beta_{n-1,1}}$ may be replaced by an isomorphic summand of the form $\mathfrak{r}_{\beta_{n-1,d}}$ for some $1\leq d\leq n-2$, and a summand of the form $\mathfrak{s}_{\alpha_{n-1,1}}$ may be replaced by an isomorphic summand $\mathfrak{s}_{\alpha_{n-1,d}}$ for some $1\leq d\leq n-3$. This implies by induction that (5.8.1) is a surjection, and implies for each $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$ with $\nu\neq\tau$ (resp. $\nu=\tau$) that $\mathfrak{r}_{\left\{\nu,\tau\right\}}$ is a simple Lie algebra acting irreducibly on $S^{\left\{\nu,\tau\right\}}$ (resp. that $\mathfrak{r}_{\left\{\nu,\pm\right\}}$ is a simple acting irreducibly on $S^{\left\{\nu,\pm\right\}}$).
Let $\mathfrak{h}=\rho_{\left\{\lambda,\mu\right\}}(\mathfrak{r}^{\prime})$. If $\left\{\nu,\nu\right\}\prec\left\{\lambda,\mu\right\}$ with $\nu=\nu^{*}$ and $|\nu|$ odd, then $\mathfrak{r}_{\left\{\nu,\pm\right\}}\cong\mathfrak{sl}\{\nu,\pm\}$ by induction, and it follows that $S^{\left\{\nu,+\right\}}$ and $S^{\left\{\nu,-\right\}}$ are dual but not isomorphic as $\mathfrak{h}$-modules because $\dim(S^{\left\{\nu,\pm\right\}})=\frac{1}{2}\cdot\binom{n-1}{|\nu|}\cdot f^{\nu}\cdot f^{\nu}\geq\frac{1}{2}\cdot(n-1)>2$, and the natural module for $\mathfrak{sl}(k)$ is self-dual only if $k=2$. Now $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ such that the restriction of $S^{\left\{\lambda,\mu\right\}}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. Since $\mathfrak{h}\subseteq\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{sl}\{\lambda,\mu\}$, then $\operatorname{rk}(\mathfrak{h})\leq\operatorname{rk}(\mathfrak{d}_{\left\{\lambda,\mu\right\}})\leq\dim(S^{\left\{\lambda,\mu\right\}})$. Our goal now is to show that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{\left\{\lambda,\mu\right\}})$. This will imply first by [Marin:2007, Lemme 15] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}$ is a simple Lie algebra, and then by [Marin:2007, Lemme 13] that $\mathfrak{d}_{\left\{\lambda,\mu\right\}}=\mathfrak{sl}\{\lambda,\mu\}$. We break the remainder of the proof into the following cases:
@@
If $\tau\prec\lambda$, then $\dim(S^{\left\{\tau,[1]\right\}})=(n-1)\cdot f^{\tau}\geq 5$. And since $\left\{\lambda,[1]\right\}$ is not an arm and leg, then $\lambda$ is not either [n1][n-1] or $1^{n-1}]$, and hence $\dim(S^{\left\{\lambda,\emptyset\right\}})=f^{\lambda}\geq(n-1)-1\geq 4$. This implies that $2\cdot\operatorname{rk}(\mathfrak{h})>\dim(S^{\left\{\lambda,[1]\right\}})$, as desired.
@@
Suppose there exists an arm and leg $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$, and all $\left\{\eta,\sigma\right\}\prec\left\{\lambda,\mu\right\}$ are proper. Replacing $\left\{\lambda,\mu\right\}$ with $\left\{\lambda^{*},\mu^{*}\right\}$ if necessary, we may assume that $\left\{\lambda,\mu\right\}=\left\{[n-1-i,1],[1^{i}]\right\}$ for some $2\leq i\leq n-2$. If $\left\{\nu,\tau\right\}\neq\left\{\nu^{*},\tau^{*}\right\}$ for all $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$, then $2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\mu\right\}})>0$ as in Section 5.4.4, so suppose $\left\{\nu,\tau\right\}=\left\{\nu^{*},\tau^{*}\right\}$ for some $\left\{\nu,\tau\right\}\prec\left\{\lambda,\mu\right\}$. Then $n=2i+1$, $\left\{\lambda,\mu\right\}=\left\{[i,1],[1^{i}]\right\}$, and as $\mathfrak{r}$-modules,
@@
Let $n\geq 6$, and suppose Theorem 5.1.1 is true for the value n1n-1. Then part (9) of Theorem 5.1.1 is true for the value $n$, i.e.,
@@
It follows from Lemma 5.9.1 that the right side of (5.9.1) is a summand of the right side of (4.3.17) for the value n1n-1. This implies by induction that (5.9.1) is a surjection, and implies for each $\nu\prec\lambda$ that $\mathfrak{r}_{\left\{\nu,\lambda\right\}}\cong\mathfrak{sl}\{\nu,\lambda\}$. Let $\mathfrak{h}=\rho_{\left\{\lambda,\pm\right\}}(\mathfrak{r}^{\prime})$. Then $\mathfrak{h}\cong\rho(\mathfrak{r}^{\prime})$ is a semisimple subalgebra of $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ such that the restriction of $S^{\left\{\lambda,\pm\right\}}$ to each simple ideal of $\mathfrak{h}$ admits a simple factor of multiplicity one. One has